Linus jogs at x kph to overtake Charlie who jogs at 6 kph. If Charlie started two hours earlier, how many hours will it take before Linus overtakes Charlie?
(Translating Word Phrases into Mathematical Expression)

Answer by stanbon(75887)

Linus jogs at x kph to overtake Charlie who jogs at 6 kph. If Charlie started two hours earlier, how many hours will it take before Linus overtakes Charlie?
----
Linus DATA:
rate = x km/h ; time = t hrs ; distance = xt km
---------------------
Charlie DATA:
rate = 6 km/h ; time = (t-2) hrs ; distance = 6(t-2) km
------
Equation:
Solve for "t"
distance = distance
xt = 6(t-2)
xt = 6t-12
6t-xt = 12
(6-x)t = 12
t = 12/(6-x) hrs
=========================
Cheers,
Stan H.